What volume is occupied by 28.2 g of argon gas at a pressure of 1.10 atm and a temperature of 326 K ?
Answer #1

Molar mass of Ar = 40 g.mol-1.

Thus, 28.2 g of Ar = (28.2 / 40) mol = 0.705 mol

From ideal gas equation we know,

PV = nRT

nRT Or, V = 0.705mol x 0.0821 L.atm.K-1.mol-1 x 326K 1.10atm = 17.1541

So, the answer is 17.154 L.